OpenStudy (anonymous):

Verify the following trigonometric identities: ((sen^2α + cos^2α)^2+(sen^2α-cos^2α)^2=2 Can someone please help?

OpenStudy (anonymous):

what is sen?

OpenStudy (anonymous):

sin* sorry

OpenStudy (anonymous):

lol.ok u know the formula of sin2x and cos2x

OpenStudy (anonymous):

sin2x=2sinxcosx

OpenStudy (anonymous):

cos2x=cos^2 x-sin^2 x

OpenStudy (ness9630):

Have you tried distributing?

OpenStudy (anonymous):

I tried to put everything on one side = 0, but I'm not sure if what I'm doing is right??

OpenStudy (anonymous):

Hello? I'm reaally not sure of what to do here

OpenStudy (crashonce):

recall the identity: sin^2A+cos^2A = 1

OpenStudy (crashonce):

that should sovle the first bit

OpenStudy (crashonce):

using ths same identity, convert the cos^2A (in the second bracket) to sin^2A then solve

OpenStudy (mathteacher1729):

This statement is not true: \[ ((\sin^2(α) + \cos^2(α))^2+(\sin^2(α)-\cos^2(α))^2=2 \] Did you mean to write a plus instead of a minus in the 2nd set of parenthesis? \[ ((\sin^2(α) + \cos^2(α))^2+(\sin^2(α)\color{red}{+}\cos^2(α))^2=2 \]

OpenStudy (mathteacher1729):

I don't know what happened to the code... :( I think \(\LaTeX\) does not like copy/pasting symbols like alpha directly into the code and replaces them with strings of ???s.

OpenStudy (anonymous):

\[[\sin ^{2}\alpha + \cos ^{2}\alpha ]^{2} + [\sin ^{2}\alpha - \cos ^{2}\alpha]^{2} = 2\]

OpenStudy (anonymous):

You should take account that sin^2= 1- cos^2 or cos^2= 1 - sin^2 \[[1-\cos ^{2}\alpha+\cos ^{2}\alpha]^{2} + [1-\cos ^{2}\alpha+\cos ^{2}\alpha]^{2}=2\]

OpenStudy (anonymous):

I am sorry. for my first equation, it should have been \[[\sin ^{2}\alpha + \cos ^{2}\alpha]^{2} + [\sin ^{2}\alpha + \cos ^{2}\alpha]^{2}\]

OpenStudy (anonymous):

carrying on from where I was ... \[1^{2} + 1^{2} = 2\]

OpenStudy (anonymous):

and the identity is proved.